# Minimal Algorithmic Information Loss Methods for Dimension Reduction, Feature Selection and Network Sparsification

### v11

@@ 2.1 Calculating information difference values
All methods employed in this article are founded on the information difference among an object’s elements. Specifically, we focus on the contribution of each element to the system’s overall information content, such as the impact of nodes or links in a network context. Formally, this involves assessing the impact of local perturbations (changes made to individual elements) on the algorithmic information content of the entire object, compared to its prior state. These methods are grounded in algorithmic/causal perturbation analysis, a key component of algorithmic information dynamics (AID) [13, 60, 40]. In the specific context of this article, our approach entails the perturbation of system elements through the removal of components, followed by measuring and ranking the effects on the system’s algorithmic information content. This dynamic approach to studying and calculating the potential changes in an object for evaluating the contribution of its components was initially introduced in [61, 64, 72, 71].
@@
which quantifies the information difference of each edge $e_{i}$ in the context of edge deletion. Notice that this difference could be defined for perturbations on the graph’s vertices, but here, we will restrict our attention to perturbations on edges. In the present article, as introduced in the next Section 2.2, we focus on the empirical investigation of destructive perturbations (i.e., the edge deletion problem) to perform network reduction while minimising information loss. However, in the Sup. Mat. we also theoretically investigate the edge-insertion case as an analogous inverse case of edge deletion. Future research will be fruitful for empirically correlating the present results in Section 4 (which hold for destructive perturbations) with those for constructive perturbations.
@@ 2.2 An algorithm for minimal information loss selection
Let $G$ be a graph, and $\left|E\left(G\right)\right|$ denote its number of edges. MILS seeks to reduce $G$ to a graph on $N<\left|E\left(G\right)\right|$ edges so that the loss of information is minimized. Ideally, one would do so by calculating the information differences between $G$ and $G\backslash S$, where $\left|S\right|\geq 1$, and then finding the subset $F\subset E(G)$ such that $I(G,F)\leq I(G,S)$ for all non-empty proper subsets of edges $S\subset E(G)$, repeating this task with $G\coloneqq G\backslash F$ until the target size is reached (i.e., when $|E(G)|=N$). This algorithmic procedure’s time complexity is clearly in $O(exp)$ because of how it performs searches and deletions across all subsets of edges (see also Fig. 2), but significant improvements to this bound are still possible as shown by Algorithm 1 (see also Algorithms $2$ and $3$ in the Sup. Mat.).
A more efficient but suboptimal version of such an exponential-time MILS iterates over single elements (in the case, edges) or singletons. Algorithm $2$ in the Sup. Mat. performs such a sequential deletions of edges, and then removes the edge that contributes less to the information content of the graph before moving to the next step. When $e$ is such that $\left|I(G,e)\right|$ is minimal, where $I(G,e)$ denotes $I(G,\{e\})$ (i.e., the information difference from deleting only a single edge), we call it a neutral information edge [72] because it is an edge that contributes the least to the information content of $G$ (in particular, it minimizes information loss or the introduction of spurious information into the network according to the information difference when removed from the original network). The existence of such edges whose information is maximally neutral (or redundant) is demonstrated in Theorem $3.8$ in the Sup. Mat., where it is shown that there are infinite families of graphs for which each member has at least one maximally neutral information edge $e_{i}$ (in the edge deletion case), where $\left|I(G,e_{i})\right|=\left|C(G)-C(G\backslash e_{i})\right|=\mathbf{O}(1)$. This is the opposite case of a maximal information edge, i.e., a single edge whose deletion changes the most the information content of the entire $G$, case which refers to incompressible graphs (see Section 3 and Fig. 1 discussed below). Thus, although suboptimal in general, our application of more efficient versions of MILS based on single edges is capable of grasping the full range between minimal and maximal information contribution to the whole network. See Section 3 for the information difference error of our efficient algorithms in comparison to the optimal exponential-time MILS.
@@
It is also important to notice that the decision tree diagram (i.e., the one for the minimisation of the information difference between the parent and the child vertex) in Fig. 2 can be subsumed into a probability tree diagram. As shown in Section $2.8.1$ in the Sup. Mat., each path from the root to a leaf has an associate joint probability of a sequence of independent events, and each event corresponds to a destructive perturbation on a subset of edges. The rooted path highlighted in the color blue in the Fig. 2 illustrates the path to the final leaf vertex (i.e., the vertex representing the outcome of Algorithm 1) that maximises the (algorithmic) probability of edge deletions, while each edge in this path corresponds to an iteration in [Algorithm 1, Step 13]. As demonstrated in Section $2.8.1$ in the Sup. Mat., which in turn is a result that evinces the generality of our methods, the MILS tree diagram in Fig. 2 shows that a decision process of minimisation of information differences implies a maximization of (algorithmic) probabilities both locally for each arrow and globally for each path from the root to a leaf. This holds theoretically independent of the choice of the computation model, programming language, or any arbitrary (computable) probability (semi-)measure that one aims at maximising (whether or not one assumes independence between parent vertices and subsequent child events in such a measure). See Section $2.8.1$ in the Sup. Mat. for a thorough explanation of the tree diagram and its mathematical properties.
@@ 3 Theory
These theoretical results pave the way for understanding why MILS algorithms are optimal with respect to their chosen edge-deletion scheme, respectively, once one fixes the arbitrarily chosen method capable of approximating the algorithmic complexity values in the asymptotic limit. That is, the algorithms return smaller networks whose respective edge count reductions are done so that the information difference between the original network and the smaller one is minimised according to those optimal upper and lower bounds described in the above paragraphs. This holds not only for the ideal EXPTIME version of MILS mentioned in Section 2.2, whose iterations cover the power set of the set $E(G)$ of present edges, but our results show that switching to a summarization procedure based on successive single-edge deletions (like in Algorithm 1) is sound. In particular, as discussed in this section, the latter case is grounded on the fact we proved that any maximum error does not dependent on the particular ideal subset of edges to be deleted, but only on the sheer size of this subset (once the worst-case logarithmic contribution of the network size becomes fixed). This means that in the theoretical limit with unbounded resources, Algorithm 1 can never misestimate the minimisation of information difference by an error larger than a linear factor of the number of single-edge perturbations that should otherwise be ideally deleted. Even if one compares the Algorithm 1’s results with the ones obtained from the best deletion scheme (i.e., the one in EXPTIME), the maximum information difference error (when choosing a scheme distinct from the ideal exponential-time MILS) increases linearly with a logarithmic term of the network size as the number of single-edge perturbations increases. Thus, this error is far from the quadratic dependence on the network size (see Corollary $3.3$ in the Sup. Mat.) that one might expect from a more simplistic method of encoding subsets of edges, as discussed to the above.
@@ 4 Results
4.1 A comparison between MILS and other data dimensionality reduction methods
@@
Figs. 34567, and 8 illustrate the effectiveness of the Minimal Information Loss Selection (MILS) method in retaining crucial local and global characteristics of synthetic and natural networks, encompassing varied types and topologies. These figures demonstrate MILS’s performance, typically surpassing leading graph sparsification algorithms. The study involved well-known networks from [49], such as genetic regulatory, protein, power grid, and social networks. MILS was applied to these networks for comparative analysis with transitive reduction [14] and spectral sparsification [58], two powerful sparsification methods. The former seeks to minimise edges while maintaining the same reachability within a directed graph, while the latter, thoroughly introduced in [19], aims to reduce network dimensions based on spectral similarity of graph Laplacians, thus preserving significant properties through the adjacency matrix Laplacian spectrum. Specifically, Figs. 3 and 4 showcase the performance of MILS in preserving the degree distribution and edge betweenness distribution in a typical synthetically (recursively) generated Erdős-Rényi (ER) random graph with low edge density, comparing its results against both random edge deletion and spectral sparsification methods.
Figs. 6, 7 and 8 illustrate how MILS outperforms spectral sparsification at preserving edge betweenness, and degree and eigenvector centralities, respectively.
4.2 An application of MILS to image lossy compression
5 Conclusion
In this paper, we have introduced Minimal Information Loss Selection (MILS), a novel agnostic approach for data summarisation that is particularly effective in reducing data dimensionality while preserving essential features such as degree distribution, clustering coefficient, edge betweenness, and centralities. As studied in Section 4.1, MILS demonstrates comparable or superior performance to existing network reduction algorithms on different classes and types of networks. Moreover, it extends to lossy compression of images and bi-dimensional data as shown in Section 4.2. It operates on the principle of local perturbation analysis and algorithmic information dynamics, offering robustness against varying methods of approximating algorithmic complexity. The method presented in this article optimally preserves the algorithmic information content (which might be affected by the local perturbations) whenever an optimal method is used to approximate algorithmic complexity. This guarantees that its performance can only increase as more computational resources are made available, which differs from other methods based on statistics for which theoretical limitations are known to hold. Nevertheless, the empirical application of MILS only scales linearly with the worst-case time complexity of the chosen method for approximating algorithmic complexity.
Funding
FSA acknowledges support from the São Paulo Research Foundation (FAPESP), grants $2023$/$05593$-$1$ and $2021$/$14501$-$8$. NAK is supported by Horizon $2020$ EU grant (MultipleMS, $733161$).
Appendix A Figures
@@
Figure 5: MILS mean clustering coefficient preservation against two other sophisticated graph sparsification methods based on graph spectral and transitive reduction techniques on biological, electric and social networks taken from [49]. The transitive method does not allow selection of edges to be deleted, and in some cases it either fails to significantly reduce the network edge density if no cycles are present (such as, generally, in electric and genetic networks) and/or takes the clustering coefficient to 0 (e.g. for protein networks) if cycles are only local. Comparisons with other methods are unnecessary because they destroy local or global properties by design, such as clustering coefficients for the spanning tree algorithm.

### v12

@@ 3.1 Calculating information difference values
All methods employed in this article are based on information difference among an object’s elements. Specifically, we focus on the contribution of each element to the system’s overall information content, such as the impact of nodes or links in a network context. Formally, this involves assessing the impact of local perturbations (changes made to individual elements) on the algorithmic information content of the entire object, compared to its prior state. These methods are grounded in algorithmic/causal perturbation analysis, a key component of algorithmic information dynamics (AID) [31, 32]. In this contribution, our approach involves the perturbation of the elements of the system through the removal of components, followed by measuring and ranking the effects on the algorithmic information content of the system. This dynamic approach to studying and calculating the potential changes in an object to assess the contribution of its components was initially introduced in [38, 37].
@@
which quantifies the information difference of each edge $e_{i}$ in the context of edge deletion. Notice that this difference could be defined for perturbations on the graph’s vertices, but here, we will restrict our attention to perturbations on edges. In the present article, as introduced in the next Section 3.2, we focus on the empirical investigation of destructive perturbations (i.e., the edge deletion problem) to perform network reduction while minimizing information loss. However, in the Sup. Mat. we also theoretically investigate the edge-insertion case as an analogous inverse case of edge deletion. Future research will be fruitful for empirically correlating the present results in Section 5 (which hold for destructive perturbations) with those for constructive perturbations.
@@ 3.2 An algorithm for minimal information loss selection
Let $G$ be a graph, and $\left|E\left(G\right)\right|$ denote its number of edges. MILS seeks to reduce $G$ to a graph on $N<\left|E\left(G\right)\right|$ edges so that the loss of information is minimised. Ideally, one would do so by calculating the information differences between $G$ and $G\backslash S$, where $\left|S\right|\geq 1$, and then finding the subset $F\subset E(G)$ such that $I(G,F)\leq I(G,S)$ for all non-empty proper subsets of edges $S\subset E(G)$, repeating this task with $G\coloneqq G\backslash F$ until the target size is reached (i.e., when $|E(G)|=N$). This algorithmic procedure’s time complexity is clearly in $O(exp)$ because of how it performs searches and deletions across all subsets of edges (see also Fig. 2), but significant improvements to this bound are still possible as shown by Algorithm 1 (see also Algorithms $2$ and $3$ in the Sup. Mat.).
A more efficient but suboptimal version of such an exponential-time MILS iterates over single elements (in the case, edges) or singletons. Algorithm $2$ in the Sup. Mat. performs such a sequential deletions of edges, and then removes the edge that contributes less to the information content of the graph before moving to the next step. When $e$ is such that $\left|I(G,e)\right|$ is minimal, where $I(G,e)$ denotes $I(G,\{e\})$ (i.e., the information difference from deleting only a single edge), we call it a neutral information edge [24] because it is an edge that contributes the least to the information content of $G$ (in particular, it minimises information loss or the introduction of spurious information into the network according to the information difference when removed from the original network). The existence of such edges whose information is maximally neutral (or redundant) is demonstrated in Theorem $3.8$ in the Sup. Mat., where it is shown that there are infinite families of graphs for which each member has at least one maximally neutral information edge $e_{i}$ (in the edge deletion case), where $\left|I(G,e_{i})\right|=\left|C(G)-C(G\backslash e_{i})\right|=\mathbf{O}(1)$. This is the opposite case of a maximal information edge, i.e., a single edge whose deletion changes the most the information content of the entire $G$, case which refers to incompressible graphs (see Section 4 and Fig. 1 discussed below). Thus, although suboptimal in general, our application of more efficient versions of MILS based on single edges is capable of grasping the full range between minimal and maximal information contribution to the whole network. See Section 4 for the information difference error of our efficient algorithms in comparison to the optimal exponential-time MILS.
@@
It is also important to notice that the decision tree diagram (i.e., the one for the minimisation of the information difference between the parent and the child vertex) in Fig. 2 can be subsumed into a probability tree diagram. As shown in Section $2.8.1$ in the Sup. Mat., each path from the root to a leaf has an associate joint probability of a sequence of independent events, and each event corresponds to a destructive perturbation on a subset of edges. The rooted path highlighted in the color blue in the Fig. 2 illustrates the path to the final leaf vertex (i.e., the vertex representing the outcome of Algorithm 1) that maximises the (algorithmic) probability of edge deletions, while each edge in this path corresponds to an iteration in [Algorithm 1, Step 13]. As demonstrated in Section $2.8.1$ in the Sup. Mat., which in turn is a result that evinces the generality of our methods, the MILS tree diagram in Fig. 2 shows that a decision process of minimisation of information differences implies a maximization of (algorithmic) probabilities both locally for each arrow and globally for each path from the root to a leaf. This holds theoretically independent of the choice of the computation model, programming language, or any arbitrary (computable) probability (semi-)measure that one aims at maximizing (whether or not one assumes independence between parent vertices and subsequent child events in such a measure). See Section $2.8.1$ in the Sup. Mat. for a thorough explanation of the tree diagram and its mathematical properties.
@@ 4 Theory
These theoretical results pave the way for understanding why MILS algorithms are optimal with respect to their chosen edge-deletion scheme, respectively, once one fixes the arbitrarily chosen method capable of approximating the algorithmic complexity values in the asymptotic limit. That is, the algorithms return smaller networks whose respective edge count reductions are done so that the information difference between the original network and the smaller one is minimised according to those optimal upper and lower bounds described in the above paragraphs. This holds not only for the ideal EXPTIME version of MILS mentioned in Section 3.2, whose iterations cover the power set of the set $E(G)$ of present edges, but our results show that switching to a summarisation procedure based on successive single-edge deletions (like in Algorithm 1) is sound. In particular, as discussed in this section, the latter case is grounded on the fact we proved that any maximum error does not dependent on the particular ideal subset of edges to be deleted, but only on the sheer size of this subset (once the worst-case logarithmic contribution of the network size becomes fixed). This means that in the theoretical limit with unbounded resources, Algorithm 1 can never misestimate the minimisation of information difference by an error larger than a linear factor of the number of single-edge perturbations that should otherwise be ideally deleted. Even if one compares the Algorithm 1’s results with the ones obtained from the best deletion scheme (i.e., the one in EXPTIME), the maximum information difference error (when choosing a scheme distinct from the ideal exponential-time MILS) increases linearly with a logarithmic term of the network size as the number of single-edge perturbations increases. Thus, this error is far from the quadratic dependence on the network size (see Corollary $3.3$ in the Sup. Mat.) that one might expect from a more simplistic method of encoding subsets of edges, as discussed to the above.
@@
4.1 How and Why MILS works
The MILS process is a greedy algorithm that begins with patch decomposition, where the input data–an image, matrix, or graph–is divided into smaller regions or patches. These patches have their algorithmic complexity estimated using BDM, which quantifies how difficult these patches are to generate as the result of a short-length program (or many of them, according to Algorithmic Probability). During the deletion phase, MILS applies perturbations such as edge deletions in graphs or bit-flipping in images while maximizing preservation of the size of the shortest programs that generate the data. Instead of relying on conventional downsampling techniques, which often distort structural information, MILS strategically removes elements in a way that minimally affects BDM (its algorithmic complexity estimation), thereby attempting to preserve the original dataset’s causal, recursive or algorithmic structure. If after removing an element the complexity of the 2D array does not change, with high probability the removed element is not relevant to the generative process itself and can therefore be safely removed as compared to other more relevant.
Then coarse-graining is performed by ensuring that the algorithmic complexity of each patch remains nearly constant even after deletion, meaning that the essential features of the data are retained despite the data dimensionality reduction. The remaining patches are then stitched together in a way that maintains their original relationships. If a region of the dataset is highly compressible and exhibits low BDM, it can tolerate more deletions without significantly affecting its algorithmic information content connected to its causal content. Conversely, highly complex regions with high BDM are preserved more carefully, as they contain intricate patterns that are difficult to reconstruct from limited data.
Ultimately, MILS does not remove data arbitrarily but rather follows an algorithmic-complexity-preservation strategy, selecting deletions that minimise changes in the generative complexity of the data. By treating the dataset as a set of algorithmic programs and ensuring that deletions do not substantially alter the shortest program needed for reconstruction, MILS effectively reduces dimensionality while maintaining structural integrity. This makes it highly valuable for applications in data compression, robust machine learning, and high-dimensional data analysis, where retaining essential information is critical.
@@ 5 Results
5.1 Experimental Validation and Benchmarking of MILS for Dimensionality Reduction in Machine Learning Pipelines
As indicated, MILS method offers a novel approach by leveraging principles of algorithmic complexity to achieve superior compression with minimal degradation of accuracy. MILS is particularly relevant in quantisation scenarios where preserving information per bit is crucial, such as low-power edge computing, compressed neural networks, and large-scale retrieval-augmented systems (RAGs). By incorporating MILS into ML pipelines, models can maintain higher accuracy per bit of compressed data, making it an attractive alternative to conventional dimension reduction methods.
To evaluate the efficiency of MILS, we conduct a benchmark study focusing on classification accuracy performance per bit of compressed image. This metric provides insight into how well different dimension reduction methods retain information relative to their storage and computational cost. The MILS method is compared against several established techniques across different quantisation levels to determine its effectiveness in preserving predictive performance while minimizing information loss.
We utilize the MNIST dataset from the ‘sklearn‘ library as the benchmark dataset for this study. It consists of 1797 8x8 images of digits, encompassing 17 features in each image (17 grayscale tones). There are about 180 samples per digit class in this dataset and it is suitable for quick benchmarking due to its relatively small size. Given its widespread use in evaluating ML techniques, MNIST serves as an ideal benchmark to test the efficacy of MILS in a controlled setting.
Three levels of quantisation are considered for our analysis and were obtained after using the ‘gfloat’ Python package:
  • 1-bit binary representation (used by MILS)
  • 4-bit floating point representation (OCP E2M1)
  • 8-bit floating point representation (OCP E4M3)
  • 16-bit floating point representation (Brain Floating Point bfloat16)
Besides, MILS is compared against three classes of dimension reduction techniques:
  • Matrix decomposition-based methods: Principal Component Analysis (PCA) and Non-Negative Matrix Factorisation (NMF).
  • Random projection-based methods: Sparse Random Projection (SPJ) and Gaussian Random Projection (GRJ).
  • Spectral methods: Kernel PCA with radial basis function (KPCAR) and polynomial kernels (KPCAP). The link of Kernel PCA with spectral methods was previously established in [40].
To properly compare compression techniques of images, we need to consider their reconstructed counterparts to assess whether information loss has occurred. Thus, rather than using the compressed representations directly, we analyze the reconstructed images to determine if the compression workflow led to a degradation of relevant information. Thus, the following Machine Learning pipeline was implemented to perform the benchmarks:
  • Create an image compression step, where different models were used to compress the original 8x8 images. The output of this step is also a 8x8 image (the one coming from reconstruction).
  • From the compressed-reconstructed version of the 8x8 image, train a Support Vector Classifier to classify the dataset according to the MNIST classes (digits from 0 to 9). Since we will perform model selection, a 5-fold StratifiedShuffleSplit cross-validation (CV) is performed with test size equal to 20% of input samples, ensuring that class balance is retained during training and that samples are shuffled. For simplicity, no hyperparameter tuning will be performed, thus a validation dataset is not needed. For the CV results assessment, the mean balanced accuracy metric was chosen (the average of recall obtained on each class).
This Pipeline was repeated for varying compression rates, hereby governed by the number of components used in traditional techniques or the number of edges kept in the MILS workflow. We chose the following values for these: 1,2,4,8,16,32 and 64, being the latter equivalent to using the full uncompressed image as input.
Besides the pyMILS implementation described in the Supplementary Material, we used another implementation which considers Algorithm 2 (where each edge is visited to check for candidate deletion cases). This MILS implementation is shown in Listing 1.
Listing 1: MILS core function
bdm = BDM(ndim=2,partition=PartitionCorrelated)
pe = PerturbationExperiment(bdm, metric=’bdm’)
def MILS(X,N):
X_red = X.copy()
idx = np.argwhere(X_red)
while idx.shape[0]>N:
pe.set_data(X_red)
delta_bdm = pe.run(idx)
b = idx[np.argmax(-np.abs(delta_bdm))]
X_red[*b] = 0
idx = np.argwhere(X_red)
return X_red
Such implementation relies on the ‘pyBDM’ Python package and its implementation of a Perturbation Experiment (which changes elements in a 2D array and calculates the resulting change in BDM). By using this function, we can define the compressor-reconstruction step for the MILS algorithm as shown in Listing 2.
Listing 2: MILS step
def MILS_step(N,input_series):
s1,s2 = input_series.shape
mils_v = Parallel(n_jobs=32)(delayed(MILS)(*x) for x in [[z,N] for z in input_series.reshape(s1,8,8)])
return np.array(mils_v).reshape(s1,s2)
In short, MILS_step performs the deletion of edges for the whole input dataset passed to the function. Several input samples are processed in parallel to enhance the speed of the transformation. It is important to notice that we consider that the compressed and reconstructed versions of an image after MILS are the same, following the rationale that changes are carried out in place. This compression-reconstruction step is then incorporated to the full Pipeline as presented in Listing 3.
Listing 3: MILS Pipeline
start = time.monotonic()
perf_MILS=[]
for nel in [2**k for k in range(7)]:
transformerMILS = FunctionTransformer(partial(MILS_step,nel))
sss = StratifiedShuffleSplit(n_splits=5, test_size=0.2, random_state=0)
pipelineMILS = Pipeline([(’transformer’, transformerMILS),
(’classifier’, SVC(gamma=’auto’))])
scr = cross_val_score(pipelineMILS, (X>7).astype(int), y, n_jobs=-1, cv = sss,scoring=’balanced_accuracy’)
perf_MILS.append([nel,scr])
end = time.monotonic()
general_time.append(["MILS",end-start])
It can be seen that prior to the application of MILS, the data is binarized by simply attributing 1 to feature values greater than 7 and 0 otherwise. By running the Pipeline in Listing 3, the performance of MILS can be assessed. To illustrate how other methods were implemented, Listing 4 presents the equivalent Pipeline for PCA reduction technique.
Listing 4: PCA Pipeline
def PCA_step(N,input_series):
mdl = Pipeline([(’transformer’, StandardScaler()),(’pca’, PCA(n_components=N))])
mdl = mdl.fit(input_series)
mdl_outp = mdl.named_steps["pca"].inverse_transform(mdl.transform(input_series))
return mdl_outp
perf_PCAs_quant=[]
for fmt,tm in zip([format_info_ocp_e2m1,format_info_ocp_e4m3,format_info_bfloat16],["PCA_e2m1","PCA_e4m3","PCA_bfloat16"]):
start = time.monotonic()
perf_PCA_q=[]
for nel in [2**k for k in range(7)]:
sss = StratifiedShuffleSplit(n_splits=5, test_size=0.2, random_state=0)
transformerPCA = FunctionTransformer(partial(PCA_step,nel))
pipelinePCA = Pipeline([(’transformer’, transformerPCA),
(’quantizer’,FunctionTransformer( lambda x: round_ndarray(fmt, x,sat=True))),
(’classifier’, SVC(gamma=’auto’))])
scrPCA = cross_val_score(pipelinePCA, round_ndarray(fmt, X,sat=True), y, n_jobs=-1, cv = sss,scoring=’balanced_accuracy’)
perf_PCA_q.append([nel,scrPCA])
perf_PCAs_quant.append(perf_PCA_q)
end = time.monotonic()
general_time.append([tm,end-start])
Listing 4 presents the implementation of a PCA_step, similar to MILS_step, where the compression and reconstruction is carried out. Differently from the case of MILS, where only a binary quantisation is considered, for all the other methods the 4-bit,8-bit and 16-bit quantisations are used. Before fitting each model, no additional preprocessing techniques were applied except for the PCA case, where a StandardScaler() (subtracts the mean and divides by the standard deviation of each feature) was applied. Also, all the simulations considered the ‘scikit-learn’ implementation of the algorithms.
Figure 3 presents the performance assessment results, which demonstrate that MILS consistently delivers higher accuracy per bit across all quantisation scenarios in the MNIST classification task. The mean balanced accuracy per bit is calculated as the mean balanced accuracy of the 5-fold CV divided by $64\times q$, where $q$ is the quantization level of the compressed-reconstructed figure.
Figure 3: Performance assessment (mean accuracy per bit of compressed image) of different compression algorithms, including MILS and other traditional approaches. MILS substantially outperforms all the considered algorithms due to its highly compressed binary setting.
Notably, as the number of components/edges is reduced, the performance gap between MILS and competing methods narrows, but MILS maintains a lead in all tested cases and quantisation levels. We attribute the superiority of MILS to its ability to optimally preserve essential structural and algorithmic complexity features of the data, leading to better generalisation in classification tasks. This is done while keeping the memory footprint to a minimum, since only binary values are used.
Figure 4, however, shows that the MILS implementation used leads to high computational cost (hereby expressed as compute time) when compared to other methods. The compute time for MILS is at least an order of magnitude higher than conventional methods, largely due to its reliance on pre-computed algorithmic complexity measures that are only available for small (4x4) patches. This limitation may be mitigated through the optimisation of MILS-specific implementation, including more efficient parallel computing strategies. Also, it is crucial to expand the precomputed Complexity-Theoretic Measure (CTM) database for image patches larger than 4×4, reducing lookup overhead.
Figure 4: Compute time in seconds for all the methods considered in this particular ML pipeline. MILS is less memory intensive at the cost of increasing computing times.
Overall, MILS presents a highly promising alternative to traditional statistical Machine Learning dimensionality reduction algorithms, particularly in quantized ML scenarios. Its ability to leverage algorithmic complexity for superior information preservation makes it especially valuable for applications in Large Language Models (LLMs) and retrieval-augmented systems (RAGs) because MILS relies exclusively on binary operations, with potential for hardware acceleration and similarity search optimisation. Future work should focus on refining its computational efficiency to facilitate broader adoption in practical ML applications.
5.2 Comparisons with other data dimensionality reduction methods
@@
Figs. 56789, and 10 illustrate the effectiveness of the Minimal Information Loss Selection (MILS) method in retaining crucial local and global characteristics of synthetic and natural networks, encompassing varied types and topologies. These figures demonstrate MILS’s performance, typically surpassing leading graph sparsification algorithms. The study involved well-known networks from [41], such as genetic regulatory, protein, power grid, and social networks. MILS was applied to these networks for comparative analysis with transitive reduction [1] and spectral sparsification [9], two powerful sparsification methods. The former seeks to minimise edges while maintaining the same reachability within a directed graph, while the latter, thoroughly introduced in [2], aims to reduce network dimensions based on spectral similarity of graph Laplacians, thus preserving significant properties through the adjacency matrix Laplacian spectrum.
@@
Specifically, Figs. 5 and 6 showcase the performance of MILS in preserving the degree distribution and edge betweenness distribution in a typical synthetically (recursively) generated Erdős-Rényi (ER) random graph with low edge density, comparing its results against both random edge deletion and spectral sparsification methods.
Figure 7: MILS mean clustering coefficient preservation against two sophisticated graph sparsification methods based on graph spectral and transitive reduction techniques on biological, electric and social networks taken from [41]. The transitive method does not allow selection of edges to be deleted, and in some cases it either fails to significantly reduce the network edge density if no cycles are present (such as, generally, in electric and genetic networks) and/or takes the clustering coefficient to 0 (e.g. for protein networks) if cycles are only local. Comparisons with other methods are unnecessary because they destroy local or global properties by design, such as clustering coefficients for the spanning tree algorithm.
Figs. 8, 9 and 10 illustrate how MILS outperforms spectral sparsification at preserving edge betweenness, and degree and eigenvector centralities, respectively.